Liberty Services is now at the end of the final year of a project. The equipment orginally cost $22,500, of which 80% has been depreciated. The firm can sell the used equipment today for $6,000, and its tax rate is 25%. What is the equipment's after-tax salvage value for use in a capital budgeting analysis?
Answer #1

book value of asset = cost - depreciation

= 22500 - 22500 * 80%

= 4500

tax on sale = (6000 - 4500) * 0.25

= 375

after tax salvage value = sale price  - tax on sale

= 6000 - 375

= 5625